The time to fracture for a vitreous silica-glass fiber at +50 ∘C is 104 s. What will be the time to fracture at room temperature (25 ∘C) ? Assume the activation energy is 78.6 kJ/mol.
Answer #1


According to clausits-clapeyson equation: log che non e(t ta) KI = ? Ti = 25+273 = 298k, Te = 50+273 + 323k log( rolete) = 2.